# Interpretable and Equation-Free Response Theory for Complex Systems

### v1

@@ IV.1.2 Linear Response - Correlations
Similarly, a good agreement is found when considering the sensitivities with respect to $\epsilon_{1}$ and $\epsilon_{2}$. In all cases it is apparent that a little fraction of the signal is lost when performing the coarse graining. Nonetheless, the good performance obtained even in the case of the 3-state system indicates the effectiveness of the reduced order modelling strategy. Using the decomposition presented in Eq. III, we are able to separate the change in the correlation function in four components, which are associated with fundamentally different dynamical processes. Two terms come from the sensitivity in the expectation value of the two observables we are considering. An additional term - indicated by $\delta^{(1)}_{b,\epsilon}$ comes from the change in the expectation value of the lagged product of the observable due exclusively to the change in the measure (where instead the evolution occurs according to the unperturbed dynamics). The most interesting term is undoubtedly $\delta^{(1)}_{a,\epsilon}$, which measures the impact of the change in the dynamics occurring up to the considered time lag. Indeed, this terms vanishes as $\tau\rightarrow 0$ and is, as already observed in Lucarini and Wouters (2017), a specific element of response formulas for correlations. The interplay between the two terms $\delta^{(1)}_{a,\epsilon}$ and $\delta^{(1)}_{b,\epsilon}$ is nontrivial.
IV.2 Nonlinear Response
Whilst explicit formulas for nonlinear Green’s functions exist Ruelle (1998b); Lucarini (2008), their numerical implementation is extremely challenging because of the convoluted structure of differential operators acting at different times. The nonlinear Green functions can be formally seen as Volterra kernels Orcioni (2014) and can in principle be constructed using neural networks Wray and Green (1994).
@@ IV.2 Nonlinear Response
Figure 4: Second order Green’s functions. Note that decay rates are controlled by a suitable combination of the two eigenvalues of the Koopman operator. Results obtained using RMM. See Eqs. IV.2-30.
Figure 5: Second order Green’s functions. Note that decay rates are controlled by a suitable combination of the two eigenvalues of the Koopman operator. Results obtained using the Ulam discretization. See Eqs. IV.2-30.
@@ V Conclusions
Constructing accurate and efficient response operators for complex systems is a problem of both theoretical and practical relevance across multiple fields in quantitative sciences Marconi et al. (2008); Baiesi and Maes (2013); Ruelle (2009); Lucarini et al. (2020). In the case of systems possessing smooth invariant measures, it is possible to resort to non-standard formulations of the FDT to recover such response operators. By combining such formalism with Koopmanism Budišić et al. (2012), one gains the important property of interpretability, as it is possible to decompose the Green’s functions of interest into a sum of terms, each associated with a specific mode of variability of the system Santos Gutiérrez and Lucarini (2022); Zagli et al. (2024). The use of Koopmanism is instrumental for establishing response formulas valid also in the case the stochastic component of a system includes jump processes Chekroun et al. (2024).
@@
A key problem in the use of response theory is that one usually needs full knowledge of the evolution equations in order to construct the Green functions and translate them in usable objects at algorithmic level. The latter task is extremely daunting especially when one deals with systems obeying deterministic evolution laws Wang (2013); Chandramoorthy and Wang (2020); Ni (2023); Ni and Tong (2024). Recently, it has been proposed to derive response operators by deploying fairly sophisticated machine learning methods based on generative score model Giorgini et al. (2024). Such a strategy, despite its great potential, is not a silver bullet for cracking the problem of constructing response operators for high dimensional models because of the need to train properly and extensively the surrogate model and issues with out-of-sample performance.
@@
It is indeed possible to devise workarounds to derive Green’s functions even in the case of extremely high dimensional systems by performing suitably defined perturbation experiments, as done in the case of climate models, where the direct evaluation of response operators seems an insurmountable task Ragone et al. (2016); Lucarini et al. (2017); Lembo et al. (2020). The flip side is that one can indeed construct useful and accurate black-box-like objects that translate forcings into predicted response, but lack ability to disentangle and possibly organize hierarchically the impact of the multiple ongoing physical processes. Hence, the level of interpretability of the response operators, despite their skills, can again be disappointing.
@@
Here we have developed a rather comprehensive set of response formulas - linear and nonlinear, for observables and for correlations, for static as well as for general time dependent perturbations - for Markov chains possessing a unique ergodic invariant measure, which are of primary importance for dynamical systems theory and statistical mechanics as a whole. See Appendix A and Appendix B for the most general formulas. Appendix C presents a linear response formulas for the total entropy production of a Markov chain undergoing time-dependent perturbation. Hopefully these results can be useful for advancing our understanding of the sensitivity of Markov chains to perturbations and, in particular, of the response near criticality, associated with the closure of the spectral gap of the unperturbed transfer operator Lucarini (2016); Santos Gutiérrez and Lucarini (2020). We will focus on this specific and extremely important problem in a separate study.
It is tempting to explore the case of absorbing Markov chains, which describe processes where there is a hole (or a trap) in the reference state space, so that the process is eventually killed Collet et al. (2013). In many cases it is possible to define and prove the existence of quasi-stationary and quasi-ergodic measures, which are constructed by adapting the usual notions to this specific case where one needs to take into account of the continuous leaks occurring in the state space Castro et al. (2024). Since the existence of such measures require, roughly speaking, the presence of a spectral gap of the transfer operator, it seems interesting to explore whether response formulas could be developed also for such Markov chains.
@@
We have thoroughly explored the efficacy of our strategy on a simple 2D gradient flow forced by additive and diagonal gaussian noise. The system we have investigated features three competing metastable states. We have shown that even considering a very severely reduced discrete representation of the system, we are able obtain high-quality information on its response to perturbations and to associated the response to specific modes of unperturbed variability.
@@
The framework we have in mind foresees the use of suitable methods of reduction of complexity of a system via MSM before the response theory developed here is used. MSM is very effective in creating surrogate representation of a possibly multiscale, many particle system in a moderate number of states Husic and Pande (2018). All one needs is a reference dataset plus few extra datasets produced with a slightly perturbed dynamics, thus allowing to obtain an estimate of the unperturbed Perron-Frobenius operator and of its perturbation in the basis defined by the MSM. Indeed, one can then study the response of the system directly at the desired coarse grained level, bypassing the need to look into all the intricacies of the system in the original resolution. The wide availability of software tools facilitating the construction of MSM Senne et al. (2012); Scherer et al. (2015); Harrigan et al. (2017), the wide range of areas of applications for MSM Husic and Pande (2018); Pande et al. (2010), recently extended to also to climate applications Springer et al. (2024), and the growing evidence of the efficacy of MSM in capturing the correlation properties of the full system Suárez et al. (2021) supports the strategy proposed here.
@@
A final comment follows. In Lucarini and Chekroun (2023b, 2024) we explored the fact that response theory provides solid foundations for a key statisical analysis method used in climate science, the optimal fingerprinting method for detection and attribution of climate change Hasselmann (1997); Allen and Tett (1999); Hegerl and Zwiers (2011); Hannart et al. (2014). Also the results presented in this contribution seem to provide some clarification to a widely used statistical analysis method used, in this case, in many signal processing applications. In Eqs. 7-8 we have shown that the Koopman operator-based expansion of the linear Green’s functions allows to decomposite it in a sum of exponential terms with weighting factors that depend on the chosen observables and on the applied forcings, whilst the exponential decay rates depend exclusively on the properties of the unperturbed system. Indeed, this functional representation points directly to the popular Prony method, which aims at representing the multivariate response collected at discrete times of a general system to instantaneous perturbations as a weighted sum of exponentials Hua and Sarkar (1990); Park and Schapery (1999); Kunis et al. (2016); Fernández Rodríguez et al. (2018). Usually, the number of exponentials one needs to use is a free and uncertain metaparameter of the statistical method. Our approach provides an interpretation of such a metaparameter, which corresponds to the number of discrete states (plus one) we consider in a hypothetical Markov state representation of the system. The well-known uncertainty in defining the optimal value for the metaparameter in the presence of strong noise and/or limited amount of available data can be linked to the difficulty in constructing an accurate Markov model in such conditions.
@@ Appendix C Entropy Production
where we have adapted Seifert’s results Seifert (2005) to the case of discrete time dynamics. We want to expand the previous expression up to first order in $\epsilon$. We obtain after lengthy but straightforward calculations:

### v2

@@ IV.1.2 Linear Response - Correlations
Similarly, a good agreement is found when considering the sensitivities with respect to $\epsilon_{1}$ and $\epsilon_{2}$. In all cases it is apparent that a little fraction of the signal is lost when performing the coarse graining. Nonetheless, the good performance obtained even in the case of the 3-state system indicates the effectiveness of the reduced order modelling strategy. Using the decomposition presented in Eq. III, we are able to separate the change in the correlation function in four components, which are associated with fundamentally different dynamical processes. Two terms come from the sensitivity in the expectation value of the two observables we are considering. An additional term - indicated by $\delta^{(1)}_{b,\epsilon}$ comes from the change in the expectation value of the lagged product of the observable due exclusively to the change in the measure (where instead the evolution occurs according to the unperturbed dynamics). The most interesting term is undoubtedly $\delta^{(1)}_{a,\epsilon}$, which measures the impact of the change in the dynamics occurring up to the considered time lag. Indeed, this terms vanishes as $\tau\rightarrow 0$ and is, as already observed in [48], a specific element of response formulas for correlations. The interplay between the two terms $\delta^{(1)}_{a,\epsilon}$ and $\delta^{(1)}_{b,\epsilon}$ is nontrivial.
IV.1.3 Nonlinear Response
Whilst explicit formulas for nonlinear Green’s functions exist [13, 14], their numerical implementation is extremely challenging because of the convoluted structure of differential operators acting at different times. The nonlinear Green functions can be formally seen as Volterra kernels [63] and can in principle be constructed using neural networks [64].
@@ IV.1.3 Nonlinear Response
Figure 4: Second order Green’s functions. Note that decay rates are controlled by a suitable combination of the two eigenvalues of the Koopman operator. Results obtained using RMM. See Eqs. IV.1.3-30.
Figure 5: Second order Green’s functions. Note that decay rates are controlled by a suitable combination of the two subdominant eigenvalues of the Koopman operator. Results obtained using the Ulam discretization. See Eqs. IV.1.3-30.
@@ V Conclusions
Constructing accurate and efficient response operators for complex systems is a problem of both theoretical and practical relevance across multiple fields in quantitative sciences [6, 2, 9, 65]. In the case of systems possessing smooth invariant measures, it is possible to resort to non-standard formulations of the FDT to recover such response operators. By combining such formalism with Koopmanism [19], one gains the important property of interpretability, as it is possible to decompose the Green’s functions of interest into a sum of terms, each associated with a specific mode of variability of the system [17, 24]. The use of Koopmanism is instrumental for establishing response formulas valid also in the case the stochastic component of a system includes jump processes [25].
@@
A key problem in the use of response theory is that one usually needs full knowledge of the evolution equations in order to construct the Green functions and translate them in usable objects at algorithmic level. The latter task is extremely daunting especially when one deals with systems obeying deterministic evolution laws [66, 67, 68, 69]. Recently, it has been proposed to derive response operators by deploying fairly sophisticated machine learning methods based on generative score model [70]. Such a strategy, despite its great potential, is not a silver bullet for cracking the problem of constructing response operators for high dimensional models because of the need to train properly and extensively the surrogate model and issues with out-of-sample performance.
@@
It is indeed possible to devise workarounds to derive Green’s functions even in the case of extremely high dimensional systems by performing suitably defined perturbation experiments, as done in the case of climate models, where the direct evaluation of response operators seems an insurmountable task [71, 72, 73]. The flip side is that one can indeed construct useful and accurate black-box-like objects that translate forcings into predicted response, but lack ability to disentangle and possibly organize hierarchically the impact of the multiple ongoing physical processes. Hence, the level of interpretability of the response operators, despite their skills, can again be disappointing.
V.1 Towards Equation-Free and Interpretable Response Formulas
@@
Here we have developed a rather comprehensive set of response formulas - linear and nonlinear, for observables and for correlations, for static as well as for general time dependent perturbations - for Markov chains possessing a unique ergodic invariant measure, which are of primary importance for dynamical systems theory and statistical mechanics as a whole. See Appendix A and Appendix B for the general formulas. Appendix C presents a linear response formulas for the total entropy production of a Markov chain undergoing time-dependent perturbation. Hopefully these results can be useful for advancing our understanding of the sensitivity of Markov chains to perturbations and, in particular, of the response near criticality, associated with the closure of the spectral gap of the unperturbed transfer operator [41, 42]. We will focus on this specific and extremely important problem in a separate study. As for future investigations, it is also tempting to explore the case of absorbing Markov chains, which describe processes where there is a hole (or a trap) in the reference state space, so that the process is eventually killed [74]. In many cases it is possible to define and prove the existence of quasi-stationary and quasi-ergodic measures, which are constructed by adapting the usual notions to this specific case where one needs to take into account of the continuous leaks occurring in the state space [75]. Since the existence of such measures require, roughly speaking, the presence of a spectral gap of the transfer operator, it seems interesting to explore whether response formulas could be developed also for such Markov chains.
@@
We have thoroughly explored the efficacy of our strategy on a simple 2D gradient flow forced by additive and diagonal gaussian noise. The system we have investigated features three competing metastable states. We have been able to construct linear and nonlinear response operators that have allow us to define sensitivity and explore response to time-dependent perturbations for observables as well as for correlations of the system. We have shown that even considering a very severely reduced discrete representation of the system, we are able obtain high-quality information on its response to perturbations and to associated the response to specific modes of unperturbed variability.
@@
The framework we have in mind foresees the use of suitable methods of reduction of complexity of a system via MSM before the response theory developed here is used. MSM is very effective in creating a surrogate representation of a possibly multiscale, many particle system in a moderate number of states [39]. All one needs is a reference dataset plus few extra datasets produced with a slightly perturbed dynamics, thus allowing to obtain an estimate of the unperturbed Perron-Frobenius operator and of its perturbation in the basis defined by the MSM. Indeed, one can then study the response of the system directly at the desired coarse grained level, bypassing the need to look into all the intricacies of the system in the original resolution. The wide availability of software tools facilitating the construction of MSM [76, 77, 78], the wide range of areas of applications for MSM [39, 37], recently extended to also to climate applications [79], and the growing evidence of the efficacy of MSM in capturing the correlation properties of the full system [80] supports the strategy proposed here.
Markov State Modelling often relies on constructing the finite state space by taking advantage of optimal Voronoï tessellations of the phase space of the system. Hence, we see a clear link between the pipeline discussed above and the combination of response theory and Koopmanism - see our recent attempt in [24] - where the operation of constructing the approximate Koopman operator is performed using the Multiplicative DMD [26], which, in turn, relies on using the characteristic functions of the cells of a Voronoï tessellation as Koopman dictionary. The relationship between these two reduced order modelling approaches will be explored in the future.
V.2 A Comment of the Prony Method
@@
A final comment follows. In [55, 81] we explored the fact that response theory provides solid foundations for a key statistical analysis method used in climate science, the optimal fingerprinting method for detection and attribution of climate change [82, 83, 84, 85]. Also the results presented in this contribution seem to provide some clarification to a widely used statistical analysis method used, in this case, in many signal processing applications. In Eqs. 7-8 we have shown that the Koopman operator-based expansion of the linear Green’s functions allows to decomposite it in a sum of exponential terms with weighting factors that depend on the chosen observables and on the applied forcings, whilst the exponential decay rates depend exclusively on the properties of the unperturbed system. Indeed, this functional representation points directly to the popular Prony method, which aims at representing the multivariate response collected at discrete times of a general system to instantaneous perturbations as a weighted sum of exponentials [86, 87, 88, 89]. Usually, the number of exponentials one needs to use is a free and uncertain metaparameter of the statistical method. Our approach provides an interpretation of such a metaparameter, which corresponds to the number of discrete states (plus one) we consider in a hypothetical Markov state representation of the system. The well-known uncertainty in defining the optimal value for the metaparameter in the presence of strong noise and/or limited amount of available data can be linked to the difficulty in constructing an accurate Markov model in such conditions.
@@ Appendix C Entropy Production
where we have adapted Seifert’s results [47] to the case of discrete time dynamics. We want to expand the previous expression up to first order in $\epsilon$. We obtain after lengthy but straightforward calculations: